So we have a normal distribution of test scores. Mean score is 120; std deviation is 20.
What is the z-value of a test score of 160??

What is the probability of selecting a test score from the group.... a test score between 120 and 160??


Answer by stanbon(75887)

So we have a normal distribution of test scores.
Mean score is 120; std deviation is 20.
What is the z-value of a test score of 160??
z(160) = (160-120)/20 = 2
----------------------------------
What is the probability of selecting a test score from the group.... a test score between 120 and 160??
z(120) = (120-120)/20 = 0
--------------------------
P(120 < x < 160) = P(0 < z < 2) = 0.4773
=======================
Cheers,
Stan H.